OpenStudy (anonymous):

find ABC ( 7x -19 ) ( 4x + 2 ) ?

OpenStudy (anonymous):

@LieutenantGeneral @dan815 @I_Need_Help_With_Home @shinalcantara

OpenStudy (lieutenantgeneral):

( 7x -19 ) ( 4x + 2 ) = 28 x^2 - 62 x - 38 x = -1/2 x = 19/7